# Detecting Harmful Memes with Decoupled Understanding and Guided CoT Reasoning

### v1

@@ Appendix D Case Study
We present examples of correctly classified memes, along with the corresponding LLM-generated reasoning outputs from Figure 4 to Figure 9. As can be seen, for these correct predictions, the rationales provided by Qwen2.5-14B and Mistral-12B are comprehensive and accurate, effectively capturing the directly related sociocultural contexts, commonsense knowledge, and the specific reasons why the memes are regarded harmful. This further verifies that the inherent understanding of small-scale LLMs regarding social norms, moral standards, and the perception of harmful content is generally aligned with human preferences.
Figure 3 shows the distribution of different error types across the three datasets of FHM, Harm-P and PrideMM. Notably, we find varying degrees of annotation errors across these three datasets, highlighting the potential bias in human annotation. Examples of such potential annotation errors are shown in Figure 13. Examples of detailed reasoning outputs of error cases are demonstrated in Figure 10, Figure 11 and Figure 12.
Figure 3: Distribution of error types within the thirty randomly sampled error cases from each of FHM, PrideMM, and Harm-P.
@@
Figure 4: Example LLM reasoning outputs for correctly detected harmful memes in FHM.
@@
Figure 5: Example LLM reasoning outputs for correctly detected harmful memes in Harm-C.
@@
Figure 6: Example LLM reasoning outputs for correctly detected harmful memes in Harm-P.
@@
Figure 7: Example LLM reasoning outputs for correctly detected harmful memes in MultiOFF.
@@
Figure 8: Example LLM reasoning outputs for correctly detected harmful memes in PrideMM.
@@
Figure 9: Example LLM reasoning outputs for correctly detected harmful memes in MAMI.
Figure 10: Error Analysis on FHM.
Figure 11: Error Analysis on PrideMM.
Figure 12: Error Analysis on Harm-P.
Figure 13: Examples of potential annotation errors in FHM, PrideMM and Harm-P.
Figure 14: Our proposed High-fidelity Meme2Text pipeline.

### v2

@@ Appendix G CASE STUDY
We present examples of correctly classified memes, along with the corresponding LLM-generated reasoning outputs from Figure 7 to Figure 12. As can be seen, for these correct predictions, the rationales provided by Qwen2.5-14B and Mistral-12B are comprehensive and accurate, effectively capturing the directly related sociocultural contexts, commonsense knowledge, and the specific reasons why the memes are regarded harmful. This further verifies that the inherent understanding of lightweight LLMs regarding social norms, moral standards, and the perception of harmful content is generally aligned with human preferences. We find varying degrees of annotation errors across the datasets, highlighting the potential bias in human annotation. Examples of such potential annotation errors are shown in Figure 16. Examples of detailed reasoning outputs of error cases are demonstrated in Figure 13, Figure 14 and Figure 15.
@@
Figure 7. Example LLM reasoning outputs for correctly detected harmful memes in FHM.
@@
Figure 8. Example LLM reasoning outputs for correctly detected harmful memes in Harm-C.
@@
Figure 9. Example LLM reasoning outputs for correctly detected harmful memes in Harm-P.
@@
Figure 10. Example LLM reasoning outputs for correctly detected harmful memes in MultiOFF.
@@
Figure 11. Example LLM reasoning outputs for correctly detected harmful memes in PrideMM.
@@
Figure 12. Example LLM reasoning outputs for correctly detected harmful memes in MAMI.
Figure 13. Error Analysis on FHM.
Figure 14. Error Analysis on PrideMM.
Figure 15. Error Analysis on Harm-P.
Figure 16. Examples of potential annotation errors in FHM, PrideMM and Harm-P.